Multiply the binomials.
(3/4 a^2 + 3b^2 ) and 4( a^2 - 2/3 b^2 )

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(3/4 a^2 + 3b^2 ) and 4( a^2 - 2/3 b^2 = (3/4 a^2 + 3b^2 ) * 4( a^2 - 2/3 b^2 )
=(3/4 a^2 + 3b^2 ) * (4a^2 - 8/3 b^2 )
=3/4 a^2 * (4a^2 - 8/3 b^2 ) + 3b^2 * (4a^2 - 8/3 b^2 )
=3/4 a^2 * 4a^2 -3/4 a^2 * 8/3 b^2 + 3b^2 * 4a^2 - 3b^2* 8/3 b^2
=3a^4–2a^2b^2 + 12 a^2 b^2 - 8b^4
= 3a^4 + 10a^2 b^2–8b^4

Thanks